You have a solution that is composed of 0.235 moles NaCl (58.45 g/mole) and 58.0 g of H2O (d = 0.986 g/mL). What is the molarity of this solution?


1
Expert's answer
2021-12-20T02:28:03-0500

d = m/V

V = m/d

V (H2O) = 58.0 g / 0.986 g/mL = 58.8 mL (0.0588 L).

To find the molarity we divide the moles of solute by the volume of the solution expressed in L:

molarity = 0.235 moles NaCl / 0.0588 L = 4.0 M